OpenStudy (anonymous):

Which is the y-intercept of the following equation? y = 4x - 6 -6 -6/4 6 4

OpenStudy (anonymous):

it is of form y=mx+c where c is y intercept so c=-6 y-intercept =-6

OpenStudy (dls):

y=mx+c y = 4x - 6 so c=-6 :)